# 8D conformal gravity with Einstein sector, and its relation to the $Q$-curvature

### v1

@@ Abstract
We first streamline the construction of the unique six-dimensional conformal gravity action found by Lü, Pang and Pope, that admits Einstein metrics as solutions to the field equations. We then prove that there exists a unique eight-dimensional conformal gravity action that admits Einstein metrics as solutions to the field equations, and explicitly build the corresponding action. Finally, we relate these results to Branson’s $Q$-curvature and the Fefferman-Graham obstruction tensor, to conclude that on every even-dimensional space there exists a unique – up to boundary terms – conformally-invariant gravity theory that is extremised by Einstein metrics.
@@ 1 Introduction
Conformal gravity is a privileged model for an extension of Einstein’s general relativity theory, since on top of the usual diffeomorphism symmetries, it is invariant under Weyl rescalings of the metric. As a classical theory of gravity, four-dimensional conformal gravity, also called Weyl gravity, was investigated in great details by several authors, see in particular Stelle:1977ry ; Riegert:1984zz ; Mannheim:1990ya ; Klemm:1998kf ; Dzhunushaliev:1999fy ; Emparan:2001wk ; Flanagan:2006ra ; Lu:2012xu ; Lu:2015cqa . A review of the relevance of Weyl gravity throughout the last decades can be found in Scholz:2017pfo to which we refer for more references.
@@
In Section 2 we briefly review the Weyl-conformal calculus developed in Boulanger:2004eh . Then, in Section 3 we review the theories of conformal gravity in four and six dimensions, that admit an Einstein sector. In Section 4 we then discuss the notion of $Q$-curvature and illustrate it explicitly in dimensions two, four and six. Then, in Section 5 we construct the most general conformal gravity theory in eight dimensions, that admits an Einstein sector, and find that the result is unique, up to boundary terms and an additive constant proportional to the Euler characteristic. We relate this action to the eight-dimensional $Q$-curvature, to find that, up to boundary terms and an additive constant, our action coincide with the (normalised) integrated $Q$-curvature. We end the note in Section 6 with a general discussion of both the $Q$-curvature and the Fefferman-Graham obstruction tensor in arbitrary even dimension $D=2m$. We argue that there is only one conformal gravity action in even dimension, that admits an Einstein sector. It coincides with the integrated $Q$-curvature, up to normalisation, boundary terms, and additive constant proportional to the Euler characteristic $\chi(M_{2m})$ of the manifold. Correspondingly, we also discuss the uniqueness of the Fefferman-Graham obstruction tensor, which is well-known in the mathematics literature.
@@ 2 Weyl-covariant tensor calculus
where we refer to Boulanger:2004eh for the definition of the generators $\bm{\Gamma}^{\alpha}$; see also below for a few examples. The important property of the Weyl-covariant derivative $\cal D$ is that its curvature vanishes if and only if the metric is conformally flat. Explicitly, one has Boulanger:2004eh
@@ 4 Relation with Branson’s $Q$-curvature
$+(2\,(\alpha-1)-\gamma_{2})\,{\bigtriangleup\sigma}\,{\bigtriangleup\varphi}$
$+2\,(D-4+2,\alpha)\,\nabla^{\mu}\sigma\,{\bigtriangleup\nabla_{\mu}\varphi}$
$+(D-2+4\,\alpha-\beta_{1})\,\nabla^{\mu}\varphi\,{\bigtriangleup\nabla_{\mu}\sigma}$
$+(\alpha-\beta_{2})\,\varphi\,{\bigtriangleup^{2}\sigma}\,$ (26)
@@
which is divergenceless, traceless, symmetric, and identically vanishing on Einstein manifolds, as it should be. Consistently with the general structure (38), one can show (see also Anastasiou:2018mfk ; Lu:2019urr ) that ${\cal Q}_{6}$ in (42) can equivalently be written as
@@
and the variational derivative $O^{(8)}_{\mu\nu}=\frac{1}{\sqrt{|g|}}\frac{\delta{\cal Q}_{8}}{\delta g^{\mu\nu}}$ is a symmetric rank-two tensor, divergenceless, traceless, of conformal dimension 6-6, that vanishes when evaluated on an Einstein metric.
@@ 6 Discussion
The $Q$-curvature is a unique quantity, that can be defined and explicitly constructed via the ambient method of FeffermanGraham2002 ; graham2005ambient ; graham2007holographic ; see also chang2008q for a review, as well as the book juhl2009 . That the functional derivative (with respect to the inverse metric) of the integrated $Q$-curvature on an even-dimensional closed manifold is proportional to the obstruction tensor was already proven in graham2005ambient , see Theorem 1.1 therein. Moreover, in the Theorem 2.1 of the latter work, the obstruction tensor was shown to obey four defining properties. To paraphrase their Theorem 2.1, if $(M^{n},[g])$ is a conformal manifold of even dimension $n=2m\geq 4$, then there exists a natural symmetric $2$-tensor ${O}_{\mu\nu}$, called the ambient obstruction tensor, with the following properties:
@@
    Trace and divergence: ${O}_{\mu\nu}$ is trace-free and divergence-free:
    $g^{\mu\nu}{O}_{\mu\nu}=0\;,\qquad\nabla^{\mu}{O}_{\mu\nu}=0\;.$ (69)
@@
Although it is not proven in graham2005ambient that these four properties uniquely specify the obstruction tensor (up to an overall coefficient), a constructive proof can be inferred from the uniqueness of the ambient metric construction itself, explained and reviewed in fefferman2012ambient .
For an axiomatic uniqueness result for the obstruction tensor, one refers to group representation theory and the Bernstein-Gelfand-Gelfand (BGG) machinery as developed for curved geometries in CapSlovakSoucek2001 , and in particular its application Gover:2002ay to the conformal case where the obstruction tensor appears as a canonical object, unique up to normalization, corresponding to a particular irreducible representation of the conformal group. For this, it suffices to ask for a symmetric rank-two conformally-covariant tensor of weight 2n2-n that is divergenceless, tracefree, and natural – i.e., polynomial in curvature and its derivatives up to order nn\,. Moreover, the BGG argument fixes not only the representation weight and symmetries, but also the symbol type of the differential operator acting on the Weyl tensor, i.e., which combination of derivatives and contractions of the Weyl curvature the tensor linearises to, which exactly translates the second property above in Eq. (68).
Therefore, the group-representation theory results of Gover:2002ay enable one to relax Property 4 above, in the axiomatic characterization of the obstruction tensor, as it is in fact a consequence of the other three. From the perspective of the present work, this shows up in the fact that the first pointwise invariant, ${\cal I}_{1}\,$, is absolutely necessary in order for the variational derivative of the Lagrangian density to vanish on a (conformally) Einstein metric, and it is the presence of this distinguished invariant ${\cal I}_{1}$ — found for the first time in Boulanger:2004zf for $n=8$ — that relates the uniqueness of the obstruction tensor between the $Q$-curvature and the group-representation (BGG) points of view, respectively. To reiterate, the property (68) is precisely the analytic avatar of the BGG uniqueness condition. It fixes the principal symbol of the invariant, thereby selecting the unique obstruction tensor from among all possible conformally covariant, symmetric, trace-free, divergence-free rank-2 tensors of weight 2n2-n.
@@
In this work, we have reviewed and re-derived the conformal actions in dimensions 4 and 6 that admit an Einstein sector, and we have compared them with the $Q$-curvature in the corresponding dimensions. We have then explicitly built the conformal gravity action in eight dimensions that admits an Einstein sector, and have shown that it is unique, up to boundary terms and overall normalisation. We have compared it with the integral of the $Q$-curvature in eight dimensions and found that they are proportional, up to an additive constant related to the Euler characteristic of the manifold, and up to boundary terms. In arbitrary even dimension, we have argued that the conformal gravity theory that admits an Einstein sector is unique. It can be defined by the integral of the $Q$-curvature on the even-dimensional spacetime manifold.
@@
In a forthcoming work in collaboration with Giorgos Anastasiou, Ignacio J. Araya, and Rodrigo Olea, we will further study the eight-conformal gravity action we have built in this work. We plan to discuss its physical properties in relation with holography and the volume-renormalisation of 8D Einstein theory, in particular addressing the problem of the conformal completion of 8D Hilbert-Einstein action, along the lines of Anastasiou:2018mfk ; Anastasiou:2024rxe ; Anastasiou:2020mik .
@@ Acknowledgements
It was a pleasure to discuss with Giorgos Anastasiou, Ignacio J. Araya, and Rodrigo Olea. We also benefited from discussions with Xavier Bekaert, Rod Gover, Yannick Herfray, and Rodrigo Olea during the workshop Conformal higher spins, twistors and boundary calculus that took place at UMONS on 30 June – 4 July 2025, where one of us (D. R. ) presented part of the present results.

### v2

@@ Abstract
We first streamline the construction of the unique six-dimensional conformal gravity action found by Lü, Pang and Pope, that admits Einstein metrics as solutions to the field equations. We then prove that there exists a unique eight-dimensional conformal gravity action that admits Einstein metrics as solutions to the field equations, and explicitly build the corresponding action. Finally, we relate these results to Branson’s $Q$-curvature and the Fefferman-Graham obstruction tensor, to conjecture that on every even-dimensional space there exists a unique – up to boundary terms – conformally-invariant gravity theory that is extremised by Einstein metrics.
@@ 1 Introduction
Conformal gravity is a privileged model for an extension of Einstein’s general relativity theory, since on top of the usual diffeomorphism symmetries, it is invariant under Weyl rescalings of the metric. As a classical theory of gravity, four-dimensional conformal gravity, also called Weyl gravity, was investigated in great details by several authors, see in particular Stelle:1977ry ; Riegert:1984zz ; Mannheim:1990ya ; Klemm:1998kf ; Dzhunushaliev:1999fy ; Emparan:2001wk ; Flanagan:2006ra ; Oliva:2012zs ; Lu:2012xu ; Lu:2015cqa . A review of the relevance of Weyl gravity throughout the last decades can be found in Scholz:2017pfo to which we refer for more references.
@@
In Section 2 we briefly review the Weyl-conformal calculus developed in Boulanger:2004eh . Then, in Section 3 we review the theories of conformal gravity in four and six dimensions, that admit an Einstein sector. In Section 4 we then discuss the notion of $Q$-curvature and illustrate it explicitly in dimensions two, four and six. Then, in Section 5 we construct the most general conformal gravity theory in eight dimensions, that admits an Einstein sector, and find that the result is unique, up to boundary terms and an additive constant proportional to the Euler characteristic. We relate this action to the eight-dimensional $Q$-curvature, to find that, up to boundary terms and an additive constant, our action coincide with the (normalised) integrated $Q$-curvature. We end the note in Section 6 with a general discussion of both the $Q$-curvature and the Fefferman-Graham obstruction tensor in arbitrary even dimension $D=2m$. We conjecture that there is only one conformal gravity action in even dimension, that admits an Einstein sector. It coincides with the integrated $Q$-curvature, up to normalisation, boundary terms, and additive constant proportional to the Euler characteristic $\chi(M_{2m})$ of the manifold.
@@ 2 Weyl-covariant tensor calculus
where we refer to this work for the definition of the generators $\bm{\Gamma}^{\alpha}$; see also below for a few examples. The important property of the Weyl-covariant derivative $\cal D$ is that its curvature vanishes if and only if the metric is conformally flat. Explicitly, one has Boulanger:2004eh
@@ 4 Relation with Branson’s $Q$-curvature
$+(2\,(\alpha-1)-\gamma_{2})\,{\bigtriangleup\sigma}\,{\bigtriangleup\varphi}+2\,(D-4+2,\alpha)\,\nabla^{\mu}\sigma\,{\bigtriangleup\nabla_{\mu}\varphi}$
$+(D-2+4\,\alpha-\beta_{1})\,\nabla^{\mu}\varphi\,{\bigtriangleup\nabla_{\mu}\sigma}+(\alpha-\beta_{2})\,\varphi\,{\bigtriangleup^{2}\sigma}\,$
@@
which is divergenceless, traceless, symmetric, and identically vanishing on Einstein manifolds, as it should be. Consistently with the general structure (37), one can show (see also Anastasiou:2018mfk ; Lu:2019urr ; Aros:2019tjw ) that ${\cal Q}_{6}$ in (41) can equivalently be written as
@@
For a different proof that the space of local conformal invariants of weight 8-8 that are divergences on 8-manifolds is 2-dimensional, see Case:2024oih .
@@
and the variational derivative $O^{(8)}_{\mu\nu}=\frac{1}{\sqrt{|g|}}\frac{\delta{\cal Q}_{8}}{\delta g^{\mu\nu}}$ is a symmetric rank-two tensor, divergenceless, traceless, of conformal dimension 6-6, that vanishes on Einstein manifolds.
@@ 6 Discussion
The $Q$-curvature can be defined and explicitly constructed via the ambient method of FeffermanGraham2002 ; graham2005ambient ; graham2007holographic ; see also chang2008q for a review, as well as the book juhl2009 . That the functional derivative (with respect to the inverse metric) of the integrated $Q$-curvature on an even-dimensional closed manifold is proportional to the obstruction tensor was already proven in graham2005ambient , see Theorem 1.1 therein. Moreover, in the Theorem 2.1 of the latter work, the obstruction tensor was shown to obey four defining properties. To paraphrase their Theorem 2.1, if $(M^{n},[g])$ is a conformal manifold of even dimension $n=2m\geq 4$, then there exists a natural symmetric $2$-tensor ${O}_{\mu\nu}$, called the ambient obstruction tensor, with the following properties:
@@
    Trace and divergence: Traceless and divergence-free: $g^{\mu\nu}{O}_{\mu\nu}=0=\nabla^{\mu}{O}_{\mu\nu}$.
@@
As far as we understand, it is unknown whether these four properties uniquely specify the obstruction tensor; it seems likely that they do.
For possible axiomatic uniqueness results concerning the obstruction tensor, another line of investigation relies on group representation theory and the Bernstein-Gelfand-Gelfand (BGG) machinery as developed for curved geometries in CapSlovakSoucek2001 with detour complexes, and in particular its application gover2006ambient ; Branson:2006ug to the (curved) conformal case, where the obstruction tensor appears as a canonical object whose vanishing allows sequences of differential operators to be conformally invariant and to form a complex. Still, in this curved case set up, we are not aware of the way to translate in group-theoretical terms the condition that the obstruction tensor should vanish on Einstein manifolds, neither are we aware of proofs of uniqueness for the obstruction tensor.
From the perspective of the present work and in the particular dimensions six and eight, the uniqueness of the action, and hence, of the obstruction tensor, appears from the requirement that the latter should vanish on an Einstein metric. For that result, it appeared crucial that the first pointwise invariant, ${\cal I}_{1}\,$, should appear in the Lagrangian density for conformal gravity, and it is the property (67) that encodes the fact that the invariant ${\cal I}_{1}\,$ should appear in the Lagrangian density. It is tempting to conjecture that, in arbitrary even dimension, the obstruction tensor is uniquely specified by the four conditions given in graham2005ambient , and therefore that the conformal gravity action that possesses an Einstein sector is unique up to boundary terms, given by the integrated $Q$-curvature in dimension $2m$.
@@
In this work, we have reviewed and re-derived the conformal actions in dimensions 4 and 6 that admit an Einstein sector, and we have compared them with the $Q$-curvature in the corresponding dimensions. We have then explicitly built the conformal gravity action in eight dimensions that admits an Einstein sector, and have shown that it is unique, up to boundary terms and overall normalisation. We have compared it with the integral of the $Q$-curvature in eight dimensions and found that they are proportional, up to an additive constant related to the Euler characteristic of the manifold, and up to boundary terms. In arbitrary even dimension, we conjecture that the conformal gravity theory that admits an Einstein sector is unique. It can be defined by the integral of the $Q$-curvature on the even-dimensional spacetime manifold.
@@
In a forthcoming work in collaboration with Giorgos Anastasiou, Ignacio J. Araya, and Rodrigo Olea, we will further study the eight-conformal gravity action we have built in this work. We plan to discuss its physical properties in relation to holography and the volume-renormalisation of 8D Einstein theory, in particular addressing the problem of the conformal completion of 8D Hilbert-Einstein action, along the lines of Anastasiou:2018mfk ; Anastasiou:2020mik ; Anastasiou:2020zwc ; Anastasiou:2024rxe  [^3] See Jia:2021hgy for a related work..
@@ Acknowledgements
We are very grateful to Thomas Basile and Jeffrey S. Case for helping us to clarify our views on the Fefferman-Graham obstruction tensor. It was a pleasure to discuss with Giorgos Anastasiou, Ignacio J. Araya, and Rodrigo Olea. We also benefited from discussions with Xavier Bekaert, Rod Gover, Yannick Herfray, and Rodrigo Olea during the workshop Conformal higher spins, twistors and boundary calculus that took place at UMONS on 30 June – 4 July 2025, where one of us (D. R. ) presented part of the present results.